# DiVERT: Distractor Generation with Variational Errors Represented as Text for Math Multiple-choice Questions

### v1

@@ 4.4 Experimental Setup
For the prompting baselines, we use the same implementation and hyperparameters as the public code repository in Feng et al. (2024) for a fair comparison. We use the latest base LLM from the GPT-4 family, GPT-4o (as of June 13, 2024), instead of GPT-3.5, to further strengthen their performance. For DiVERT, we use MetaMath-Mistral $7$Yu et al. (2023) as our base LLM since it is one of the best-performing LLMs in the $7$B family on mathematical reasoning; we found that it outperforms other open-source LLMs with similar size on our tasks. See Supplementary Material F for detailed parameter settings. At test time, we overgenerate $N_{e}=10$ errors via $\text{LLM}^{e}$ using diverse beam search Vijayakumar et al. (2018) and $N_{d}=10$ distractors through $\text{LLM}^{d}$ using standard beam search. Among the set of $N_{e}\times N_{d}=100$ error-distractor pairs we select the top-$K\in\{3,10\}$. For DisSearch-ED CoT, we use standard beam search with $100$ beams for a fair comparison with DiVERT. DisSearch-D performs better with $10$ beams so we report its performance with $10$ beams.
@@ 5.1 Quantitative evaluation
DiVERT performs comparably or better than GPT-4o and outperforms baselines. Table 1 shows downstream distractor generation performance for all approaches on all evaluation metrics. Our variational approach, DiVERT, using an open-source base LLM with 7B parameters, performs on par with GPT-4o, especially on the robust Proportional@$10$ metric, where it outperforms GPT-4o by a wide margin. GPT-4o kNN performs best on $K=3$ evaluation metrics, which is not surprising since prior work Feng et al. (2024) found that the kNN approach exploits in-context examples in the training set that have the same underlying structure as the target MCQ, different only in named entities and numerical values. Therefore, all it needs to do is to follow patterns and generate three distractors correspondingly without understanding errors. However, it does not perform well on $K=10$ metrics since it often fails to go beyond the top three errors in MCQs that have numerous plausible errors. In contrast, DiVERT performs well by training on error labels to acquire an understanding of plausible mathematical errors. Under the same base LLM, DiVERT performs much better than baselines on almost all metrics, which highlights the importance of its variational approach, since sampling from the $\text{LLM}^{q}$ model during training encourages exploration of the error space and improves model robustness.
Errors as “chain-of-thought” improve performance. Comparing among baselines, we see that training on human-authored error explanations improves distractor generation performance. This result can easily be explained by error labels serving as valuable chain-of-thought Wei et al. (2022) supervision during training.
@@
DiVERT works even with a small number of error labels. Since soliciting error labels behind question-distractor pairs from math educators is time-consuming, we investigate DiVERT’s reliance on the amount of error label data. Figure 2 shows DiVERT’s distractor generation performance using different portions of the available error labels. We see that on the most stable metric, Prop@$10$, DiVERT performs better than baselines, especially when the majority of error labels are dropped: the performance gap is widest when less than half of the training error labels are used for training. This result further highlights the importance of the self-exploration training setup in DiVERT’s variational approach at going beyond human-authored labels and linking different error types across questions.
Ablation study. Table 2 shows results of the ablation study on the more robust @$10$ metrics. We see that the variational training process is crucial and removing it results in a large drop in performance (see Supplementary Material C). We also see that not training $q_{\phi}$ and using its fine-tuned initialized version $q_{\phi_{init}}$ also drops performance, which highlights the importance of sampling from $q_{\phi}$ and exploring the error space of math MCQs. Regularizing $q_{\phi}$ and using temperature annealing in the soft token approximation are also effective at making training more robust.
@@
Table 2: Ablation study of DiVERT on distractor generation performance on $K=10$ metrics.
@@ 5.2 Qualitative Evaluation
Table 3 shows generated errors and corresponding distractors for an MCQ stem in the test set. DiVERT generates a diverse set of plausible errors and corresponding distractors. However, GPT-4o Zero-shot CoT prioritizes error diversity and generates errors unlikely to be made by real students like “Confusing the concept of multiples with that of prime factors”. GPT-4o kNN fails to generalize on this test example; it does not generate diverse distractors and instead follows a fixed rule of generating either multiples of $12$ but not $15$, and vice versa, highlighting its pattern-following nature and dependence on good in-context examples. The best fine-tuning baseline, DisSearch-ED CoT, generates generic error descriptions like “Does not understand the term multiple”, which is not specific to the question’s context.
@@
GPT-4o kNN (feedback shown below)
This is a multiple of 12 but not of 15 … 24
This is a multiple of 15 but not of 12 … 45
This is a multiple of 12 but not of 15 … 36
@@
Failure Pattern Analysis. We now investigate failure patterns in the generated errors and corresponding distractors from DiVERT. Table 4 shows a representative example MCQ in the test set. We observe that a majority of generated errors from $p(e|s)$ are mathematically valid, with high diversity, but some are less likely among real students such as “Subtracts instead of divides”, matching the observation made in prior work Feng et al. (2024). By far, the most frequent failure pattern we observe is on the controllable distractor generation model $p(d|s,e)$, where the generated distractor is not faithful to the error. In this example, for the error “When dividing a fraction by an integer, divides the denominator by the integer”, the generated distractor is $\frac{6}{1}$ rather than the desired $\frac{6}{3}$. The reverse can also occasionally happen when a distractor is valid but the error is not. This observation highlights that although effectively learning good error representations and capable of identifying what errors can be made in a math MCQ, the biggest limitation of DiVERT is in its inability to enforce consistency in the downstream distractor generation model $p(d|s,e)$. This limitation suggests a major direction for future work, possibly by exploring the use of an error-distractor consistency penalty in the training objective.
@@ 6.1 Evaluation Setup
While automated distractor evaluation is well-defined since there is ground-truth, it is more challenging to automatically evaluate errors. Reference-based evaluation may penalize errors that faithfully reflect the mathematical error behind a distractor but are semantically different than the ground truth, or conversely, reward errors that are invalid but semantically similar to the ground truth. To address this challenge, we conduct a human evaluation to measure the quality of generated and human-authored errors. We recruit two experienced math teachers who have extensive expertise in designing math MCQs as annotators. We randomly select 20 questions from the test set on a diverse set of topics, and for each question, show annotators the ground-truth, human-authored errors, and generated errors from both DiVERT and GPT-4o. For DiVERT, we select the top 3 distractors from the $p(e|s)$ model using diverse beam search to promote error diversity. For GPT-4o, we select the errors generated during CoT distractor generation. We randomize the order of errors shown to the annotators and do not let them know how each error is generated. We instruct annotators to rate each error on a 5-point Likert scale, with 5 being the best, depending on whether an error is relevant to the question, mathematically sound, specific, conceptual, and likely to be made by a real student. See Supplementary Material I.1 for the exact instructions given to the annotators.
@@ 7 Conclusions and Future Work
Our goal in this work is to develop a system that can automatically create distractors for assessments or practice problems, and do so in an explainable way so that the distractors can be more easily verified by human educators. We hope that such systems will save educators time on content creation, allowing them to spend more resources on personal student interactions. However, there is a concern that such systems could replace human educator jobs, which is a shared concern across most domains with AI applications. We note that while there is little risk for bias in the creation of mathematical distractors, the use of textual errors introduces the possibility of generating biased text by inheriting tendencies from the base LLM. Another risk of automatically generating distractors is that a lower quality compared to human-authored content could lead to negative student learning outcomes. Because of these reasons, we recommend that errors and distractors be reviewed by experts before being deployed to real students.
Appendix A Related Work
For automated distractor generation in language learning and reading comprehension, prior works have explored ranking candidate distractors based on semantic similarity to the key and using knowledge graphs Susanti et al. (2018); Stasaski and Hearst (2017); Alsubait et al. (2014). More recent works use an end-to-end pipeline for distractor generation, which lead to longer and higher-quality distractors Qiu et al. (2020); Shuai et al. (2023); Xie et al. (2021); Gao et al. (2019), also leveraging open-source LLMs such as BERT and T5 Kalpakchi and Boye (2021); Chiang et al. (2022); Rodriguez-Torrealba et al. (2022); Qu et al. (2024); Wang et al. (2023a); these approach are similar to the baselines we use in this work. Other works prompt state-of-the-art proprietary LLMs such as ChatGPT and GPT-4 to generate distractors, with carefully crafted prompts, for a wider range of subjects including math and computer science Tran et al. (2023); Bitew et al. (2023); Feng et al. (2024). Evaluating the quality of distractors, however, remains challenging: Moore et al. (2024) proposes a series of features that can be used to evaluate the quality of distractors, although those features are mostly about surface semantics and do not apply to subjects like math that require deeper reasoning. Another line of recent work in Feng et al. (2024); Scarlatos et al. (2024) proposes to measure the quality of distractors through how likely they are going to be selected by real students, which requires training another model to predict student option selection behavior. However, most existing MCQ datasets do not come with such student response information.
Appendix B Monte Carlo Approximation of ELBO Training Objective
@@
$\mathbb{E}_{q_{\phi}(e|s,d)}$ $[\log p_{\theta_{d}}(d|s,e)]$
$-\beta D$ (${}_{\text{KL}}(q_{\phi}(e|s,d)\ ||\ p_{\theta_{e}}(e|s))$
@@
$-\beta\mathbb{E}$ [${}_{q_{\phi}(e|s,d)}[\log q_{\phi}(e|s,d)\ -\ \log p_{\theta_{e}}(e|s)]$
$\approx\sum_{\hat{e}\sim q_{\phi}(e|s,d)}$ $\log p_{\theta_{d}}(d|s,\hat{e})-\beta\log q_{\phi}(\hat{e}|s,d)$
@@ Appendix D Real-world Math MCQ Dataset Details
We divide the real-world math MCQ dataset into train-val-test splits by questions to ensure no overlap in the MCQ stem across the splits, resulting in roughly a $72\%$-$16\%$-$12\%$ split over question-distractor pairs. Table 6 shows detailed statistics of the dataset. We manually checked a random subset of the data and found no personally identifying information or offensive content.
@@ E.2 Baselines and Results
Table 7 shows error generation performance for all error-based approaches on all evaluation metrics. As a reference for the diversity of predicted errors shown, the diversity of ground-truth, human-written errors is $0.574$, $0.300$, and $0.349$, for the choice of the similarity metric as ROUGE-L F1, cosine similarity, and BERTScore F1, respectively. The finetuning baseline ErrorSearch-E imitates the ground-truth human-written error distribution and performs best on F1 performance across almost all textual similarity metrics. DisSearch-ED CoT performs best on precision performance across all textual similarity metrics. However, the same beam search decoding helping precision, leads to a drop in the performance of DisSearch-ED CoT on recall and diversity. GPT-4o Zero-shot CoT exhibits good error diversity, but as expected performs poorly on textual similarity metrics, with the zero-shot errors generated not matching the distribution of human-written errors. The $p(e|s)$ model from our variational method, DiVERT, generates errors with the highest diversity. This diversity also leads to a slight drop in overall F1 performance across textual similarity metrics. This result is not surprising since by design, during the variational training of DiVERT, the $p(e|s)$ model aligns with the entropy model $q_{\phi}(e|s,d)$, which is encouraged to generate error samples around human-written errors to learn a robust error space representation, leading to better downstream distractor generation performance.

### v2

@@ 4.4 Experimental Setup
For the prompting baselines, we use the same implementation and hyperparameters as the public code repository in Feng et al. (2024) for a fair comparison. We use the latest base LLM from the GPT-4 family, GPT-4o (as of June 13, 2024), instead of GPT-3.5, to further strengthen their performance. For DiVERT, we use MetaMath-Mistral $7$Yu et al. (2023) as our base LLM since it is one of the best-performing LLMs in the $7$B family on mathematical reasoning; we found that it outperforms other open-source LLMs with similar size on our tasks. See Supplementary Material F for detailed parameter settings. At test time, we overgenerate $N_{e}=10$ errors via $\text{LLM}^{e}$ using diverse beam search Vijayakumar et al. (2018) and $N_{d}=10$ distractors through $\text{LLM}^{d}$ using standard beam search. Among the set of $N_{e}\times N_{d}=100$ error-distractor pairs we select the top-$K\in\{3,10\}$. For DisSearch-ED CoT, we use standard beam search with $100$ beams for a fair comparison with DiVERT. DisSearch-D performs better with $10$ beams so we report its performance with $10$ beams. For the primary results in Table 1, we perform a 5-fold cross-validation, rotating the test set over the dataset, and report the average and standard deviation of metrics across methods. For all other experiments, we evaluate on a single fold to reduce training time.
@@ 5.1 Quantitative Evaluation
DiVERT performs comparably or better than GPT-4o and outperforms baselines.
Table 1 shows downstream distractor generation performance for all approaches on all evaluation metrics. Our variational approach, DiVERT, using an open-source base LLM with 7B parameters, performs on par with GPT-4o, especially on the robust Proportional@$10$ metric, where it outperforms GPT-4o by a wide margin. GPT-4o kNN performs best on $K=3$ evaluation metrics, which is not surprising since prior work Feng et al. (2024) found that the kNN approach exploits in-context examples in the training set that have the same underlying structure as the target MCQ, different only in named entities and numerical values. Therefore, all it needs to do is to follow patterns and generate three distractors correspondingly without understanding errors. However, it does not perform as well on $K=10$ metrics since it often fails to go beyond the top three errors in MCQs that have numerous plausible errors. In contrast, DiVERT performs well by training on error labels to acquire an understanding of plausible mathematical errors. Under the same base LLM, DiVERT performs better than baselines on almost all metrics, which highlights the importance of its variational approach, since sampling from the $\text{LLM}^{q}$ model during training encourages exploration of the error space and improves model robustness.
Errors as “chain-of-thought” improve performance.
Comparing among baselines, we see that training on human-authored error explanations improves distractor generation performance. This result can easily be explained by error labels serving as valuable chain-of-thought Wei et al. (2022) supervision during training.
@@
DiVERT works even with a small number of error labels.
Since soliciting error labels behind question-distractor pairs from math educators is time-consuming, we investigate DiVERT’s reliance on the amount of error label data. Figure 2 shows DiVERT’s distractor generation performance using different portions of the available error labels. We see that on the most stable metric, Prop@$10$, DiVERT performs better than baselines, especially when the majority of error labels are dropped: the performance gap is widest when less than half of the training error labels are used for training. This result further highlights the importance of the self-exploration training setup in DiVERT’s variational approach at going beyond human-authored labels and linking different error types across questions.
Ablation study.
@@
Table 2: Ablation study of DiVERT on distractor generation performance on $K=10$ metrics on a single fold.
Table 2 shows the results of the ablation study on the more robust @$10$ metrics on a single fold. We see that the variational training process is crucial and removing it results in a large drop in performance (see Supplementary Material C). We also see that not training $q_{\phi}$ and using its fine-tuned initialized version $q_{\phi_{init}}$ also drops performance, which highlights the importance of sampling from $q_{\phi}$ and exploring the error space of math MCQs. Regularizing $q_{\phi}$ and using temperature annealing in the soft token approximation are also effective at making training more robust.
@@ 5.2 Qualitative Evaluation
Table 3 shows generated errors and corresponding distractors for an MCQ stem in the test set. DiVERT and GPT-4o kNN each generate a diverse set of plausible errors and corresponding distractors. However, GPT-4o Zero-shot CoT prioritizes error diversity and generates errors unlikely to be made by real students like “Confusing the concept of multiples with that of prime factors”. The best fine-tuning baseline, DisSearch-ED CoT, generates generic error descriptions like “Does not understand the term multiple”, which is not specific to the question’s context.
@@
Ground-Truth
Believes finding the product of two numbers gives their lowest common multiple. $180$
Believes the largest number in a set of numbers is always their lowest common multiple. $15$
Confuses factors and multiples. $3$
@@
GPT-4o kNN
Confuses the product of the two numbers with the LCM. 180
Incorrectly identifies a common multiple that is not the lowest. 120
Incorrectly adds the two numbers. 27
@@
Failure Pattern Analysis.
We now investigate failure patterns in the generated errors and corresponding distractors from DiVERT. Table 4 shows a representative example MCQ in the test set. We observe that a majority of generated errors from $p(e|s)$ are mathematically valid, with high diversity, but some are less likely among real students such as “Subtracts instead of divides”, matching the observation made in prior work Feng et al. (2024). By far, the most frequent failure pattern we observe is on the controllable distractor generation model $p(d|s,e)$, where the generated distractor is not faithful/consistent to the error. In this example, for the error “When dividing a fraction by an integer, divides the denominator by the integer”, the generated distractor is $\frac{6}{1}$ rather than the consistent $\frac{6}{3}$. The reverse can also occasionally happen when a distractor is plausible but the error is not. This observation highlights that although effectively learning good error representations and capable of identifying what errors can be made in a math MCQ, the biggest limitation of DiVERT is in its inability to enforce consistency in the downstream distractor generation model $p(d|s,e)$. This limitation suggests a major direction for future work, possibly by exploring the use of an error-distractor consistency penalty in the training objective.
@@ 6.1 Evaluation Setup
While automated distractor evaluation is well-defined since there is ground-truth, it is more challenging to automatically evaluate errors. Reference-based evaluation may penalize errors that faithfully reflect the mathematical error behind a distractor but are semantically different than the ground truth, or conversely, reward errors that are invalid but semantically similar to the ground truth. To address this challenge, we conduct a human evaluation to measure the quality of generated and human-authored errors. We recruit two experienced math teachers who have extensive expertise in designing math MCQs as annotators. We randomly select $20$ questions from the test set on a diverse set of topics, and for each question, show annotators the ground-truth, human-authored errors, and generated errors from both DiVERT and GPT-4o, for a total of $180$ errors. For DiVERT, we select the top 3 distractors from the $p(e|s)$ model using diverse beam search to promote error diversity. For GPT-4o, we select the errors generated during CoT distractor generation. We randomize the order of errors shown to annotators and do not tell them how each error is generated. We instruct annotators to rate each error on a 5-point Likert scale, with 5 being the best, depending on whether an error is relevant to the question, mathematically sound, specific, conceptual, and likely to be made by a real student. See Supplementary Material I.1 for the exact instructions given to annotators.
@@
Error Distractor
Plausible error, plausible and consistent distractor.
@@ 6.2 Results
7 Related Work
For automated distractor generation in language learning and reading comprehension, prior works have explored ranking candidate distractors based on semantic similarity to the key and using knowledge graphs Susanti et al. (2018); Stasaski and Hearst (2017); Alsubait et al. (2014). More recent works use an end-to-end pipeline for distractor generation, which lead to longer and higher-quality distractors Lee et al. (2024); Qiu et al. (2020); Shuai et al. (2023); Xie et al. (2021); Gao et al. (2019), also leveraging open-source LLMs such as BERT and T5 Kalpakchi and Boye (2021); Chiang et al. (2022); Rodriguez-Torrealba et al. (2022); Qu et al. (2024); Wang et al. (2023a); these approaches are similar to the baselines we use in this work. Other works prompt state-of-the-art proprietary LLMs such as ChatGPT and GPT-4 to generate distractors, with carefully crafted prompts, for a wider range of subjects including math and computer science Tran et al. (2023); Bitew et al. (2023); Feng et al. (2024). Evaluating the quality of distractors, however, remains challenging: Moore et al. (2024) proposes a series of features that can be used to evaluate the quality of distractors, although those features are mostly about surface semantics and do not apply to subjects like math that require deeper reasoning. Another line of recent work in Feng et al. (2024); Scarlatos et al. (2024a) proposes to measure the quality of distractors through how likely they are going to be selected by real students, which requires training another model to predict student option selection behavior. However, most existing MCQ datasets do not come with such student response information.
@@ 8 Conclusions and Future Work
Acknowledgements
The authors would like to thank the annotators and Jaewook Lee, Hunter McNichols, Hasnain Heickal, and Nancy Otero for helpful discussions. We also thank the anonymous reviewers for their helpful comments. This work is partially supported by Schmidt Futures under the learning engineering virtual institute (LEVI) initiative and the NSF under grants 2118706, 2237676, and 2341948.
@@ Acknowledgements
Our goal in this work is to develop a system that can automatically create distractors for assessments or practice problems, and do so in an explainable way so that the distractors can be easily verified by human educators. We hope that such systems will save educators time on content creation, allowing them to spend more resources on personal student interactions. However, there is a concern that such systems could replace human educator jobs, which is a shared concern across most domains with AI applications. We note that while there is little risk for bias in the creation of mathematical distractors, the use of textual errors introduces the possibility of generating biased text by inheriting tendencies from the base LLM. Another risk of automatically generating distractors is that lower-quality learning content compared to human-authored content could lead to negative student learning outcomes. Because of these reasons, we recommend that generated errors and distractors be reviewed by experts before being deployed to real students.
Appendix A Monte Carlo Approximation of ELBO Training Objective
@@
$\mathbb{E}$ [${}_{q_{\phi}(e|s,d)}[\log p_{\theta_{d}}(d|s,e)]$
$-\beta D_{\text{KL}}(q_{\phi}(e|s,d)\ ||\ p_{\theta_{e}}(e|s))$
@@
$-\beta\mathbb{E}_{q_{\phi}(e|s,d)}[\log q_{\phi}(e|s,d)\ -\ \log p_{\theta_{e}}(e|s)]$
$\approx$ $\textstyle\sum_{\hat{e}\sim q_{\phi}(e|s,d)}\log p_{\theta_{d}}(d|s,\hat{e})-\beta\log q_{\phi}(\hat{e}|s,d)$
@@
Appendix B Additional Baselines
In Table 6, we show results on distractor generation for additional baselines from Feng et al. (2024), including kNN using feedback instead of errors and the rule-based method, where GPT-4o first selects errors from a pool before generating distractors. We also run the CoT baseline using the OpenAI o1 model OpenAI (2024b). To reduce costs, we collect these additional results on a single fold of the data, and show the other GPT-4o baselines and DiVERT performance on this fold for reference.
We first observe that CoT with o1 only slightly outperforms CoT with GPT-4o, improving Prop@10 by 1.62. This indicates that the long internal reasoning generated by o1 does not increase alignment with teacher-written distractors, though it may reduce arithmetic or logical errors resulting in the slight increase. We next observe that the Rule-Based method is slightly better than CoT, though still significantly worse than kNN, consistent with findings in Feng et al. (2024). Finally, we observe there is little difference when using feedback instead of errors in kNN. This result validates our choice of using errors as textual reasoning before generating distractors, and also indicates that the exact form of the reasoning may not be significant in the context of generating aligned distractors.
Model K=3 K=10
Exact@3 Partial@3 Prop@3 Exact@10 Partial@10 Prop@10
Proprietary Base LLM, OpenAI o1
OpenAI o1 Zero-shot CoT $4.17$ $72.22$ $36.11$ $25.69$ $77.78$ $52.08$
Proprietary Base LLM, GPT-4o
GPT-4o Zero-shot CoT Feng et al. (2024) $2.78$ $68.75$ $33.33$ $20.83$ $81.25$ $50.46$
GPT-4o Rule-Based Feng et al. (2024) $4.17$ $67.36$ $35.42$ $26.39$ $77.78$ $53.70$
GPT-4o kNN Feedback Feng et al. (2024) 25.69 $\underline{83.33}$ 55.09 $\underline{36.81}$ $\underline{88.89}$ $\underline{63.89}$
GPT-4o kNN Errors $\underline{23.61}$ 85.42 $\underline{52.78}$ $\underline{36.81}$ $\underline{88.89}$ $63.66$
Open-source Base LLM, MetaMath-Mistral 7B
DiVERT (ours) $13.19$ $81.25$ $46.06$ 42.36 91.67 68.75
Table 6: Single fold performance on distractor generation for additional baselines and reference methods. Best performance is in bold and second best is underlined.
@@ Appendix D Real-world Math MCQ Dataset Details
We divide the real-world math MCQ dataset into train-val-test splits by questions to ensure no overlap in the MCQ stem across the splits, resulting in roughly a $72\%$-$16\%$-$12\%$ split over question-distractor pairs. Table 7 shows detailed statistics of the dataset. Each of the $1434$ MCQs has $3$ distractors leading to $1434\cdot 3=4302$ question-distractor pairs. Among these pairs, $3601$ pairs have human annotated error descriptions; other pairs are not labeled due to the non-mathematical nature of errors, e.g., careless slipping, reading the question incorrectly, etc. There are $1231$ unique error descriptions. For the test set, we only keep questions containing human annotated error descriptions for all three associated distractors, leading to $144$ questions with $144\cdot 3=432$ question-distractor pairs. We manually checked a random subset of the data and found no personally identifying information or offensive content.
@@ E.2 Baselines and Results
Table 8 shows error generation performance for all error-based approaches on all evaluation metrics. As a reference for the diversity of predicted errors shown, the diversity of ground-truth, human-written errors is $0.574$, $0.300$, and $0.349$, for the choice of the similarity metric as ROUGE-L F1, cosine similarity, and BERTScore F1, respectively. The finetuning baseline ErrorSearch-E imitates the ground-truth human-written error distribution and performs best on recall performance across all textual similarity metrics. DisSearch-ED CoT performs best on precision performance across all textual similarity metrics. However, the same beam search decoding helping precision, leads to a drop in the performance of DisSearch-ED CoT on recall and diversity. GPT-4o Zero-shot CoT exhibits good error diversity, but as expected performs poorly on textual similarity metrics, with the zero-shot errors generated not matching the distribution of human-written errors. The $p(e|s)$ model from our variational method, DiVERT, generates errors with the highest diversity. This diversity also leads to a slight drop in overall F1 performance across textual similarity metrics. This result is not surprising since by design, during the variational training of DiVERT, the $p(e|s)$ model aligns with the entropy model $q_{\phi}(e|s,d)$, which is encouraged to generate error samples around human-written errors to learn a robust error space representation, leading to better downstream distractor generation performance.